OpenStudy (anonymous):

solve : sin squared 2x = 1

OpenStudy (anonymous):

take square root of both sides sin(2x) = 1 sin(2x) = -1 can you do it now ?

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

what is the final answer?

OpenStudy (anonymous):

2x=1 divide both side by 2 x=1/2 check: 2(1/2)=1 2/2=1 1=1 so the answer is correct

OpenStudy (anonymous):

sin(2x) = 1 -> 2x = 90 + 2*pi*n -> x = 45 + pi*n sin(2x) = -1 -> 2x = 270 + 2*pi*n -> x = 135 + pi*n -> x = 45 + (pi/2)*n